The following data is available regarding X Company utility cost in 2017: Estimated production Budgeted fixed costs Budgeted total variable costs Actual production Actual total costs 3,000 units $10,000 $20,250 2,700 units s24,654 What was the flexible budget variance for total utility costs in 2017 (assume that a positive number means a favorable variance and a negative number means an unfavorable variance)? Submit Answer Tries 0/2
Answer #1
Unit variable cost for utilities = 20250/3000= $6.75
Utility costs in Flexible budget = 10000+(2700*6.75)= $28225
Flexible budget variance for total utility costs = Costs in flexible budget-Actual total costs
Flexible budget variance for total utility costs = 28225-24654 = 3571 favorable